OpenStudy (1645323):

Voltage V is related to current I and power P by the equation V = P/I. If the power of a generator is modeled by P = t3 + 2t2 + 21t - 12. If the current of the generator is I = t + 4, write an expression that represents the voltage. Select one: a. V = t2 + 6t - 3 b. V = t2 - 6t + 4 c. V = 2t2 - 6t - 3 d. V = t + 6

OpenStudy (1645323):

@terenzreignz

OpenStudy (1645323):

?

OpenStudy (1645323):

is it too hard? my teacher ght help me with this one

OpenStudy (ybarrap):

It's tedious. But here are some details. So far, none match your options: $$ V=t^3/(t+4)+(2 t^2)/(t+4)+(21 t)/(t+4)-12/(t+4)\\ =(t-2) t-128/(t+4)+29\\ =t^2-2 t-128/(t+4)+29 $$

OpenStudy (1645323):

ok then my teacher messed it up then but thanks

OpenStudy (ybarrap):

yw